# Learning to Validate Generative Models: a Goodness-of-Fit Approach

### v1

@@ Abstract
Generative models are increasingly central to scientific workflows, yet their systematic use and interpretation require a proper understanding of their limitations through rigorous validation. Classic approaches struggle with scalability, statistical power, or interpretability when applied to high-dimensional data, making it difficult to certify the reliability of these models in realistic, high-dimensional scientific settings. Here, we propose the use of the New Physics Learning Machine (NPLM), a learning-based approach to goodness-of-fit testing inspired by the Neyman–Pearson construction, to test generative networks trained on high-dimensional scientific data. We demonstrate the performance of NPLM for validation in two benchmark cases: generative models trained on mixtures of Gaussian models with increasing dimensionality, and a public end-to-end generator for the Large Hadron Collider called FlashSim, trained on jet data, typical in the field of high-energy physics. We demonstrate that the NPLM can serve as a powerful validation method while also providing a means to diagnose sub-optimally modeled regions of the data.
@@ 1 Introduction
Generative artificial intelligence (AI) tools have drawn significant attention in scientific research, where they have demonstrated the potential to reshape the discovery workflow and have already contributed to advances in fields such as protein design, drug discovery, and materials science (see for example [NEURIPS2019_f3a4ff48, anstine2023generative, zeni2025generative]). In the path to scientific discovery, the modeling of probability densities plays a crucial role to address very complex theory calculations, parameter estimation and perform data analysis. Recent developments have introduced new highly expressive generative models and sophisticated training strategies able to capture the non-trivial structures characterizing scientific data, providing surrogate models that can be sampled efficiently. In the field of physical sciences, generative AI is providing the tools to accelerate theory calculations and numerical simulations of physical phenomena, opening the way to the statistical interpretations of very complex systems otherwise hard to handle. For instance, in high-energy physics (HEP), fast simulations could help overcome the limitations of computational and storage resources anticipated in the future high-luminosity phase of the Large Hadron Collider (LHC) [ZurbanoFernandez:2020cco]. Therefore generative networks (GNs) are being applied to a variety of complex tasks, such as LHC event generation [Butter:2022rso], calorimeter shower simulation [Krause:2024avx], anomaly detection [Kasieczka:2021xcg], likelihood modeling [Reyes-Gonzalez:2023oei, Heimel:2024drk], and more.
@@
This work is organized as follows. In Section 2, we review the practice of two-sample testing for goodness-of-fit to assess the compatibility between a given reference distribution, $p_{R}$, and a set of data from a distribution $q$. We then describe the NPLM approach. In Section 3, we present our results for two cases: the performance of normalizing flows (NFs) trained on mixtures of Gaussians with increasing dimensionality, and the performance of FlashSim [vaselli2024endtoend], an NF-based model trained on LHC jet data. In Section 4 we showcase how the method can be used as a diagnostic tool to identify mismodelled regions in the data. Finally, in Section 5, we present our conclusions and outlook.
@@ 3 Experiments
In this section, we demonstrate how to apply the NPLM method to the problem of validating and evaluating generative models. This work focuses particularly on Normalising Flows (NFs) [rezende2015variational, kobyzev2020normalizing]. These models are widely adopted in scientific applications because they provide explicit likelihoods that can be both sampled and evaluated.
Nevertheless, this approach can be applied to any generative model since the analysis methods are not specific to NFs. As outlined in Section 2, we perform two complementary sets of experiments:
@@ 3.2 Particle Jets with FlowSim
As a second use case, we study the generative model proposed in Ref. [vaselli2024endtoend]. This model, dubbed as FlashSim, is a novel approach to fast simulation of high-energy physics collision events based on Normalizing Flows and Flow Matching [lipman2023flowmatchinggenerativemodeling], which provides simulations considered in good agreement with traditional approaches based on Monte Carlo techniques, with the benefit of a significant of speed-up of orders of magnitude. The objective is to evaluate the FlashSim model, by comparing it with the traditional simulation approach based on Monte Carlo methods that we refer to as FullSim. These simulations are the standard for current Large Hadron Collider (LHC) analyses and serve as the target for replication by generative models.
@@
FlashSim aims at simulating jets, namely sprays of particles originating in proton-proton collisions observed in particle physics experiments at the LHC. [^1] See [vaselli2024endtoend] for a detailed description of the physics problem and properties of jets. Each data point in the dataset represents a jet object generated from a tt̄ process at LHC, and is described by 16 observables. The features describing each jet are divided in basic reconstructed, namely $p_{T}$, $\eta$, $\phi$, invariant mass, b-tagging discriminator and number of constituents; and extended dataset variables, namely neutral hadron fraction (nhf), charged hadron fraction (chf), neutral electromagnetic fraction (nef), charged electromagnetic fraction (cef), quark-gluon discriminator (qgd), jet identification (JetId), number of charged and neutral particles, c-tagging discriminator and number of secondary vertices (nSV). The description of these variables and how they are generated or reconstructed by a generative model is given in Section 3 of Ref. [vaselli2024endtoend].
@@
The testing procedure follows the methodology described in Section 3.1. As in the previous case, two tests are performed depending on the chosen reference model. The reference sample size is fixed to 1 million events, while the data sample size varies between 5,000 and 200,000 events. The model hyperparameters are determined according to the strategy outlined in Section 3.1, consistent with the previous benchmark, yielding $M=8000$, $\lambda=10^{-6}$, and $\sigma=7.4$. Once the NPLM model is defined, a distinct null hypothesis is estimated for each sample size. Each estimation is based on $N_{\rm toys}=200$ toy datasets. The distribution of the test statistic under the alternative hypothesis is obtained by repeating the test 40 times. Given the finite availability of FullSim data ($\approx$ 15 million events), each training is performed on a bootstrapped sample drawn from the complete FullSim dataset. Similarly, for the FlashSim model, we rely on bootstrapping from a total of approximately 4 million sampled events for the analysis.
Figure 4: Z-scores of the Flashsim validation as function of the size of the analyzed FlashSim dataset. The blue line represents FullSim as the reference case, while FlashSim as the reference case is shown in orange.
$N_{\mathcal{D}}$ FullSim as ref FlashSim as ref
@@
Table 4: Z-scores of the NPLM test for FlashSim data at varying sample size.
@@ Results
Table 4 reports the resulting Z-scores for different sample sizes. As in the previous section, the error associated to the Z-score represents the $68\%$ confidence interval around the median. In Fig. 4, we observe that the level of discrepancy between the reference model and the data increases with the sample size $N_{\mathcal{D}}$. This is expected and it is a further indication of the consistency of our testing methodology. As before, we find that using the generative model as the reference leads to lower median Z-scores. Nonetheless, the observed difference between the reference choices is now within the uncertainties. Furthermore, we estimate that in order to obtain a discrepancy lower than $1\sigma$, $3\sigma$, and $5\sigma$, the sample sizes must be approximately $10^{4}$, $4\cdot 10^{4}$, and $7\cdot 10^{4}$, respectively, when FullSim is the reference, and $10^{4}$, $5\cdot 10^{4}$, and $8\cdot 10^{4}$, respectively, when FlashSim is the reference.
@@ 5 Conclusions
The code used for the MoG study is provided in Ref. [NPLM-Flows], while the trained NF models and MoG implementations are available in Ref. [NF_MoGs_NPLM]. The code employed in the FlashSim study can be found in Ref. [NPLM-FlashSim].
@@ Appendix A NPLM model selection
where $\{\tilde{x}_{1},...,\tilde{x}_{M}\}$ are called Nyström centres and are sampled uniformly at random from the input data, with $M$ an hyperparameter. The corresponding solution can be shown to be with high probability as accurate as the exact one (see [rudi2017generalization] and references therein). Problem (8) is then solved by an approximate Newton iteration, as discussed in details in [meanti2020kernel]. Ultimately, the Falkon-based implementation of NPLM is characterised by three main hyperparameters: the number of centres $M$, the kernel width $\sigma$ and the regularisation parameter $\lambda$. These are tuned only on reference data to avoid biases toward specific anomalous features that might be present in the measurements $\mathcal{D}$. Following [Letizia:2022xbe, Grosso:2023ltd], the criteria for their selection can be summarised as:
@@
The Gaussian kernel width $\sigma$ is set at the $90\%$ quantile of the distribution of pairwise distance between reference points. Once $\sigma$ is set, we study how the median value of the NPLM test distribution under the null hypothesis varies with the number of Nystrom centers ($M$) and the regularization parameter ($\lambda$). We note that the median value of the empirical distribution of the test has a saturation region along $M$, i.e. from a certain value of $M$ the median is almost constant. We use this information to select $M$. Both values of M and $\lambda$ have to deal with the timing performance for a single toy which has to be reasonable. Thus we select the value of $\lambda$ as small as possible to preserve the kernel method complexity while keeping the execution time reasonable. Lastly, it is fundamental to consider the compatibility of the resulting empirical distribution $p(t|H_{0})$ with a $\chi^{2}$ distribution. A p-value for the Kolmorogov-Smirnov test is estimated by comparing the t distribution with the fitted $\chi^{2}$, in order to ensure the trend of the distribution $p(t|H_{0})$. This aspect is important when the final p-value is estimated.

### v2

@@ Abstract
Generative models are increasingly central to scientific workflows, yet their systematic use and interpretation require a proper understanding of their limitations through rigorous validation. Classic approaches struggle with scalability, statistical power, or interpretability when applied to high-dimensional data, making it difficult to certify the reliability of these models in realistic, high-dimensional scientific settings. Here, we propose the use of the New Physics Learning Machine (NPLM), a learning-based approach to goodness-of-fit testing inspired by the Neyman–Pearson construction, to test generative networks trained on high-dimensional scientific data. We demonstrate the performance of NPLM for validation in two benchmark cases: generative models trained on mixtures of Gaussian models with increasing dimensionality, and a public end-to-end model, known as FlowSim, developed to generate high-energy physics collision events. We demonstrate that the NPLM can serve as a powerful validation method while also providing a means to diagnose sub-optimally modeled regions of the data.
@@ 1 Introduction
Generative artificial intelligence (AI) tools have drawn significant attention in scientific research, where they have demonstrated the potential to reshape the discovery workflow and have already contributed to advances in fields such as protein design, drug discovery, and materials science (see for example [NEURIPS2019_f3a4ff48, anstine2023generative, zeni2025generative]). In the path to scientific discovery, the modeling of probability densities plays a crucial role to address very complex theory calculations, parameter estimation and perform data analysis. Recent developments have introduced new highly expressive generative models and sophisticated training strategies able to capture the non-trivial structures characterizing scientific data, providing surrogate models that can be sampled efficiently. In the field of physical sciences, generative AI is providing the tools to accelerate theory calculations and numerical simulations of physical phenomena, opening the way to the statistical interpretations of very complex systems otherwise hard to handle. For instance, in high-energy physics (HEP), fast simulations could help overcome the limitations of computational and storage resources anticipated in the future high-luminosity phase of the Large Hadron Collider (LHC) [ZurbanoFernandez:2020cco]. Therefore generative networks (GNs) are being applied to a variety of complex tasks, such as LHC event generation [Butter:2022rso], calorimeter shower simulation [Krause_2025], anomaly detection [Kasieczka:2021xcg], likelihood modeling [Reyes-Gonzalez:2023oei, Heimel:2024drk], and more.
@@
This work is organized as follows. In Section 2, we review the practice of two-sample testing for goodness-of-fit to assess the compatibility between a given reference distribution, $p_{R}$, and a set of data from a distribution $q$, and then describe the NPLM approach. In Section 3, we present our results for two cases: the performance of normalizing flows (NFs) trained on mixtures of Gaussians with increasing dimensionality, and the performance of FlowSim [vaselli2024endtoend], an NF-based model trained on simulated LHC-like jet data. In Section 4 we showcase how the method can be used as a diagnostic tool to identify mismodelled regions in the data. Finally, in Section 5, we present our conclusions and outlook.
@@ 3 Experiments
In this section, we demonstrate how to apply the NPLM method to the problem of validating and evaluating generative models. This work focuses particularly on Normalising Flows (NFs) [rezende2015variational, kobyzev2020normalizing]. These models are widely adopted in scientific applications because they provide explicit likelihoods that can be both sampled and evaluated. Nevertheless, this approach can be applied to any generative model since the analysis methods are not specific to NFs.
As outlined in Section 2, we perform two complementary sets of experiments:
@@ 3.2 Particle Jets with FlowSim
As a second use case, we study the generative model proposed in Ref. [vaselli2024endtoend]. This model, dubbed FlowSim, is a Normalizing Flow– and Flow Matching–based approach to fast simulation of high-energy collision events [lipman2023flowmatchinggenerativemodeling]. FlowSim delivers simulation quality in good agreement with traditional Monte Carlo–based methods while achieving orders-of-magnitude speed-ups. In Ref. [vaselli2024endtoend], the model is trained not on full geant4-level simulations [GEANT4:2002zbu], which would be the realistic target, but on a simplified collider dataset: events are generated with pythia8 [Bierlich:2022pfr] and passed through a lightweight, Delphes-like [deFavereau:2013fsa] toy detector that introduces smearing, flavour-dependent responses, and realistic correlations among jet kinematics, tagging variables, and constituent properties. This setup provides a controlled yet detector-like environment, with full details available in the original paper. Our objective is to evaluate FlowSim by comparing it to the traditional Monte Carlo–based simulation pipeline, referred to as FullSim, which serves as the standard for current LHC analyses and the target for generative model replication.
@@
FlowSim aims at simulating jets, namely sprays of particles originating in proton-proton collisions observed in particle physics experiments at the LHC. [^1] See [vaselli2024endtoend] for a detailed description of the physics problem and properties of jets. Each data point in the dataset represents a jet object generated from a tt̄ process at LHC, and is described by 16 observables. The features describing each jet are divided in basic reconstructed, namely $p_{T}$, $\eta$, $\phi$, invariant mass, b-tagging discriminator and number of constituents; and extended dataset variables, namely neutral hadron fraction (nhf), charged hadron fraction (chf), neutral electromagnetic fraction (nef), charged electromagnetic fraction (cef), quark-gluon discriminator (qgd), jet identification (JetId), number of charged and neutral particles, c-tagging discriminator and number of secondary vertices (nSV). An in-depth description of these variables and how they are generated or reconstructed by a generative model can be found in Ref. [vaselli2024endtoend].
@@
The testing procedure follows the methodology described in Section 3.1. As in the previous case, two tests are performed depending on the chosen reference model. The reference sample size is fixed to 1 million events, while the data sample size varies between 5,000 and 200,000 events. The model hyperparameters are determined according to the strategy outlined in Section 3, consistent with the previous benchmark, yielding $M=8000$, $\lambda=10^{-6}$, and $\sigma=7.4$. Once the NPLM model is defined, a distinct null hypothesis is estimated for each sample size. Each estimation is based on $N_{\rm toys}=200$ toy datasets. The distribution of the test statistic under the alternative hypothesis is obtained by repeating the test 40 times. Given the finite availability of FullSim data ($\approx$ 15 million events), each training is performed on a bootstrapped sample drawn from the complete FullSim dataset. Similarly, for the FlowSim model, we rely on bootstrapping from a total of approximately 4 million sampled events for the analysis.
Figure 4: Z-scores of the FlowSim validation as function of the size of the analyzed FlowSim dataset. The blue line represents FullSim as the reference case, while FlowSim as the reference case is shown in orange.
$N_{\mathcal{D}}$ FullSim as ref FlowSim as ref
@@
Table 4: Z-scores of the NPLM test for FlowSim data at varying sample size.
@@ Results
Table 4 reports the resulting Z-scores for different sample sizes. As in the previous section, the error associated to the Z-score represents the $68\%$ confidence interval around the median. In Fig. 4, we observe that the level of discrepancy between the reference model and the data increases with the sample size $N_{\mathcal{D}}$. This is expected and it is a further indication of the consistency of our testing methodology. As before, we find that using the generative model as the reference leads to lower median Z-scores. Nonetheless, the observed difference between the reference choices is now within the uncertainties. Furthermore, we estimate that in order to obtain a discrepancy lower than $1\sigma$, $3\sigma$, and $5\sigma$, the sample sizes must be approximately $10^{4}$, $4\cdot 10^{4}$, and $7\cdot 10^{4}$, respectively, when FullSim is the reference, and $10^{4}$, $5\cdot 10^{4}$, and $8\cdot 10^{4}$, respectively, when FlowSim is the reference.
@@ 5 Conclusions
The code used for the MoG study is provided in Ref. [NPLM-Flows], while the trained NF models and MoG implementations are available in Ref. [NF_MoGs_NPLM]. The code employed in the FlowSim study can be found in Ref. [NPLM-FlashSim].
@@ Appendix A NPLM model selection
where $\{\tilde{x}_{1},...,\tilde{x}_{M}\}$ are called Nyström centres and are sampled uniformly at random from the input data, with $M$ an hyperparameter. The corresponding solution can be shown to be with high probability as accurate as the exact one (see [rudi2017generalization] and references therein). Problem (8) is then solved by an approximate Newton iteration, as discussed in details in [meanti2020kernel]. Ultimately, the Falkon-based implementation of NPLM is characterised by three main hyperparameters: the number of centres $M$, the kernel width $\sigma$ and the regularisation parameter $\lambda$. These are tuned only on reference data to avoid biases toward specific anomalous features that might be present in the data of interest $\mathcal{D}$. Following [Letizia:2022xbe, Grosso:2023ltd], the criteria for their selection can be summarised as:
@@
Once $\sigma$ is set, we study how the median value of the NPLM test distribution under the null hypothesis varies with the number of Nystrom centers ($M$) and the regularization parameter ($\lambda$). We note that the median value of the empirical distribution of the test has a saturation region along $M$, i.e. from a certain value of $M$ the median is almost constant. We use this information to select $M$. Both values of M and $\lambda$ have to deal with the timing performance for a single toy which has to be reasonable. Thus we select the value of $\lambda$ as small as possible to preserve the kernel method complexity while keeping the execution time reasonable. Lastly, it is fundamental to consider the compatibility of the resulting empirical distribution $p(t|H_{0})$ with a $\chi^{2}$ distribution. A p-value for the Kolmorogov-Smirnov test is estimated by comparing the t distribution with the fitted $\chi^{2}$, in order to ensure the trend of the distribution $p(t|H_{0})$. This aspect is important when the final p-value is estimated.